OpenStudy (stickman1998):

Not sure how to solve: A sound wave is modeled with the equation. Equation in comments

OpenStudy (stickman1998):

y=1/4cos2pi/3theta

OpenStudy (stickman1998):

\[(y)=1/4\cos(2\pi/3) \theta\]